Determine the mass of a pierce of copper where the heat needed to raise the temperature of the copper from 0°C to 110 °C was 293,000 J . The specific heat capacity of copper is 376 J per kilogram per degree Celsius. Answer in whole numbers and one decimal place in kg.
Answer #1

∆Q = mc∆T

=) m = ∆Q/c∆T = 293000/376×110 = 7.08kg

=) m = 7.1kg.